OpenStudy (anonymous):

if you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x - 5y + z = 12 x + 6y - z = -4 2x - 5y - 2z = -8 possible answers: 4x + y = 8 -17y = 0 2x - 11y = 16 4x + 7y = -16 -23y + 4z = 24 -17y = 0 13y - 2z = 24 7y - 4z = 0

hartnn (hartnn):

so could u isolate x from 2nd equation?

OpenStudy (anonymous):

x= -6y + z -4

hartnn (hartnn):

correct! now put that x in 1st equation

OpenStudy (anonymous):

3(-6y + z - 4) - 5y + z =12

hartnn (hartnn):

simplify it....combine like terms

hartnn (hartnn):

-18y+3z-12-5y+z=12 -23y+4z=24 got this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think I figured it out!

hartnn (hartnn):

great! :D